Classify the following statements as true or false giving reasons for your answers.  i) General solution of the differential equations x^2(d^2y/dx^2)^(3) +y(dy/dx)^(4) +y^4 =0 must  contain four arbitrary constants.
(ii) The set of real (or complex) solutions of equation y'(x) + P(x)y(x)=Q(x) forms a
real (or complex) vector space.
(iii) sin xd^2y/dx^2 +dy/dx +y=0 ]0,π[ is linear homogeneous equation.
iv) The partial differential equation
x^2 ∂^2z/∂x^2 +2xy ∂^2z/∂x∂y +y^2 ∂^2z/∂y^2 –x^m y^n =0
is a reducible homogeneous equation.
(v) The partial differential equation u ∂u/∂x =e^y + sin x, u= u(x,y) is a quasi-linear
p.d.e.
1
Expert's answer
2020-03-02T07:26:29-0500

I) False. The order of this equation is 2, hence it should contain 2 arbitraty constants

II) True. For all "\lambda, \mu\in C" If "y_1" and "y_2" are solutions of the equation, "\lambda y_1+\mu y_2" is also the solution.

III)True. It has the form "A(x)y''+B(x)y'+C(x)y=0" therefore, it is the linear homogeneous equation of second order.

IV) True.the equation can be written in an operational form "x^2D^2+2xyDD'+y^2D'=(x^2D+y^2D')^2"

V)True. since "\sin x+e^y" is not linear